OpenStudy (anonymous):

f(x)=sin(6x+2) f'(x)=? I'm supposed to be using the chain rule, and with that I have: (cos(x))(6x+2)(6) Help please!

OpenStudy (zzr0ck3r):

no, its 6cos(6x+2)

OpenStudy (kinggeorge):

\[f'(x) = \cos(6x+2)\cdot {d \over dx} (6x+2)=6\cos(6x+2)\]

OpenStudy (anonymous):

f(x)=sin(x) f'(x)=cos(x) * x'

OpenStudy (zzr0ck3r):

this is just the chain rule, take derivative of the outside function and leave the inside the same, then multiply by the derivative of the inside function

OpenStudy (earthcitizen):

\[d/dx(\sin(6 x+2)) = 6 \cos(6 x+2)\]

OpenStudy (zzr0ck3r):

the same thing happens when we take the derivative of x^2, its really 2(x)(1) we just dont write the 1

OpenStudy (anonymous):

So then f(x)=sin(x^4) f'(x)=cos(4x^3)(x^4)

OpenStudy (kinggeorge):

\[f'(x)=\cos(x^4)4x^3\]

OpenStudy (zzr0ck3r):

no dont take the derivative on the inside ccos(x^4)(4x^3)

OpenStudy (zzr0ck3r):

leave the inside the same, just ike we leave the x alone in x^2

OpenStudy (anonymous):

Awesome thanks you wonderful people

OpenStudy (earthcitizen):

yw